OpenStudy (anonymous):

what is 4x + 2y = -20 written in slope-intercept form?

OpenStudy (solomonzelman):

y=mx+b, sompare to this and rearrange you equation to set it into the same form.

OpenStudy (anonymous):

@SolomonZelman is it y = -4x + 20?

OpenStudy (solomonzelman):

not exatcly, it is not -4x + 20 on the right.

OpenStudy (anonymous):

@SolomonZelman i think i found it! is it y = 20x - 4?

OpenStudy (solomonzelman):

nope,

OpenStudy (anonymous):

y = 2x - 20? @SolomonZelman

OpenStudy (solomonzelman):

\[4x + 2y = -20\]\[4x + 2y -4x= -20-4x\]\[2y=-4x-20\]so just y is equal to....

OpenStudy (solomonzelman):

you divide everything by 2, not just the -4x, -20 also.

OpenStudy (anonymous):

oh so then it would be y = 2x - 10? @SolomonZelman

OpenStudy (solomonzelman):

still not exactly

OpenStudy (anonymous):

or no, y = -2x - 10 because it's negative! @solomonZelman

OpenStudy (solomonzelman):

YUP!!!!

OpenStudy (solomonzelman):

(•‿•)

OpenStudy (anonymous):

@SolomonZelman ok thanks so much!

OpenStudy (solomonzelman):

Anytime